OpenStudy (idkwut):

Find the three points of intersection with the equation -x^2 = -2^x, please show step by step.

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

if the base is the same number then you just have to solve the exponents. which would show that 2=x

OpenStudy (anonymous):

do you kind of get what im saying?

OpenStudy (idkwut):

Yes. But I know that there are two other solutions, right?

OpenStudy (idkwut):

I am actually trying to find the three points of intersection.

OpenStudy (anonymous):

oh

OpenStudy (idkwut):

@myininaya @ganeshie8 Anyone?

myininaya (myininaya):

so you actually mean \[x^2=2^x ? \]

OpenStudy (idkwut):

Yea

OpenStudy (idkwut):

Acctually wait it's -x^2 = -2^x

OpenStudy (idkwut):

But unless that's another form for it

myininaya (myininaya):

so you mean it is \[-1 \cdot x^2 =-1 \cdot 2^x \]

myininaya (myininaya):

I just want to make sure you are actually looking at \[2^x=x^2\]

OpenStudy (idkwut):

ok

OpenStudy (idkwut):

My textbook just showed it as how i stated so i just left it like that

OpenStudy (idkwut):

but can you show me step by step? I know that there are at least 3 intersections

myininaya (myininaya):

well if it is the equation you said it is then we could rewrite it as \[2^x=x^2\]

OpenStudy (idkwut):

ok

myininaya (myininaya):

so are you taking calculus and using an approximation method or are you using a more of a guess method and plugging sort of method

myininaya (myininaya):

there is no straight forward algebra way to solve this

OpenStudy (idkwut):

Im taking calc so I guess it's approximation method

myininaya (myininaya):

what kind of approximation method?

OpenStudy (idkwut):

No wait, actually, can we do it the guess method way?

myininaya (myininaya):

i can only guess two solutions i can't think of a third

OpenStudy (idkwut):

which two solutions are u thinkin?

myininaya (myininaya):

that one guy said 2^2=2^2 what other integer would give us the same thing on both sides of x^2=2^x?

myininaya (myininaya):

i will give you a hint another solution is \[2^{something }\]

OpenStudy (idkwut):

4

myininaya (myininaya):

yep

myininaya (myininaya):

4^2=2^4 16=16

myininaya (myininaya):

i'm not sure if i can help you anymore though

OpenStudy (idkwut):

Ok, i just plugged this equation into wolfram and the third intersection is a -0.766, but idk how they got that

OpenStudy (idkwut):

i dont think it can solved algebraically more like a calculator approximation so i'll just list that so my teacher will know

OpenStudy (idkwut):

thank u

myininaya (myininaya):

you can use newton's method